### v1

# Practical Insights into Knowledge Distillation for Pre-Trained Models

Abstract
This research investigates the enhancement of knowledge distillation (KD) processes in pre-trained models, an emerging field in knowledge transfer with significant implications for distributed training and federated learning environments. These environments benefit from reduced communication demands and accommodate various model architectures. Despite the adoption of numerous KD approaches for transferring knowledge among pre-trained models, a comprehensive understanding of KD’s application in these scenarios is lacking. Our study conducts an extensive comparison of multiple KD techniques, including standard KD, tuned KD (via optimized temperature and weight parameters), deep mutual learning, and data partitioning KD. We assess these methods across various data distribution strategies to identify the most effective contexts for each. Through detailed examination of hyperparameter tuning, informed by extensive grid search evaluations, we pinpoint when adjustments are crucial to enhance model performance. This paper sheds light on optimal hyperparameter settings for distinct data partitioning scenarios and investigates KD’s role in improving federated learning by minimizing communication rounds and expediting the training process. By filling a notable void in current research, our findings serve as a practical framework for leveraging KD in pre-trained models within collaborative and federated learning frameworks.
I Introduction
Knowledge distillation (KD), initially conceptualized for model compression [1, 2, 3], has evolved to facilitate knowledge transfer between pre-trained models[4, 5, 6, 7]. Notably, these pre-trained models might be exposed to distinct data samples, resulting in performance variations due to various heterogeneities. One of these is statistical heterogeneity, where models are trained with data that are not independent and identically distributed (non-IID). Additionally, system heterogeneity introduces challenges, as devices vary in their available bandwidth and computational capabilities. Learning across such diverse systems and imbalanced datasets is complex yet critical.
The application of KD in distributed training and federated learning environments is increasingly recognized for its potential to mitigate communication overhead and facilitate diverse model architectures [6, 7, 8, 4, 5, 9, 10, 11]. Despite this growing interest, there remains a notable gap in understanding KD’s application, particularly in pre-trained models. Recent studies employ various KD algorithms—such as vanilla KD, deep mutual learning [4, 9], and data partitioning KD[12]—to facilitate knowledge transfer between pre-trained models. These methods’ effectiveness across varied data distribution strategies and their comparative advantages or limitations have not been systematically investigated.
Additionally, the realm of hyperparameter optimization, particularly concerning parameters such as temperature and weight, remains largely unexplored, presenting a significant opportunity for advancement in this domain [13].
This paper aims to address these gaps by conducting an in-depth study on joint distillation [13], a scenario where distributed participants engage in direct peer-to-peer knowledge exchange between pre-trained models. We present a comparative analysis of diverse KD approaches, considering various data partitioning strategies. Our exploration includes an assessment of the influence of transfer set options and a detailed investigation into hyperparameter optimization. Through extensive grid search results, we identify practical guidelines for tuning to enhance model performance. This study aims to enrich the understanding of KD techniques in pre-trained models and provide actionable insights for optimizing these methods in distributed and federated learning contexts.
This paper advances the field of knowledge distillation in distributed training systems with the following contributions:
  • We conduct a comprehensive comparative analysis of KD approaches across five data partitioning strategies and assess the impact of different transfer set options.
  • We examine hyperparameter optimization, focusing on temperature and weight, using grid search to ascertain when tuning is most beneficial and to determine optimal tuning scenarios for diverse data partitions.
  • We explore KD learning dynamics, analyzing learning-forgetting balance and criteria for effective teacher model selection across data scenarios.
  • We demonstrate the efficiency of KD in federated learning (FL), evidenced by reduced training rounds to reach desired accuracy levels.
We plan to openly release the source code associated with this study.
II Preliminaries and related work
II-A Knowledge distillation
KD is a technique initially proposed to transfer knowledge from a large, complex model (termed the teacher) to a smaller model (termed the student), with the goal of maintaining comparable performance [1, 2, 3]. The central concept is that the student model can learn more effectively by emulating the soft targets, which are the output probabilities computed using the softmax function parameterized by temperature $T$, as shown:
$p_{i}^{t}(z_{t,l},T)=\frac{\exp(z_{t,i,l}/T)}{\sum_{j}\exp(z_{t,j,l}/T)}.$ (1)
Here, $z_{t,i,l}$ and $z_{t,j,l}$ are logits for classes $i$ and $j$ from the teacher model for a specific data sample $l$. The temperature $T$ is used to control the smoothness of the probability distribution; in KD contexts, it is often set to a value greater than 1 to produce softer probability distributions [1].
Various KD methodologies utilize different loss functions and datasets, often referred to as a transfer set or proxy dataset, which can include public labeled or unlabeled datasets, or the teacher’s training data. When labeled data is available, the student model’s training combines the cross-entropy loss with the teacher’s soft targets and a distillation loss as:
$\mathcal{L}=(1-\alpha)\>\mathcal{L}_{CE}(p^{s},y)+\alpha\>\mathcal{L}_{KL}(p^{s},p^{t}).$ (2)
The weight $\alpha$ balances the cross-entropy loss $\mathcal{L}_{CE}$ and the Kullback-Leibler divergence loss $\mathcal{L}_{KL}$; $y$ represents the true labels; $p^{s}$ and $p^{t}$ denote the predicted probabilities (cf. Eq. 1) by the student and teacher models, respectively. Fig. 1 illustrates the KD pipeline.
Figure 1: Knowledge distillation pipeline.
II-B Knowledge distillation with pre-trained models
Beyond model compression, KD’s applicability includes scenarios where both the teacher and student models are pre-trained on different data subsets. This application of KD aims to consolidate the distinct knowledge each model possesses, a concept increasingly relevant in distributed training and federated learning [4, 5, 9, 6, 11, 10, 7, 12].
Our study focuses in the following approaches:
Vanilla KD
In the context of pre-trained models, Vanilla KD employs the standard KD process, (Eq. 2), where the primary focus is on transferring knowledge from one pre-trained model to another, leveraging the established KD framework.
Tuned KD
Tuned KD refers to the optimization of KD hyperparameters, particularly the temperature ($T$) and weight ($\alpha$). This approach adjusts the standard KD framework to better suit the characteristics and learning capacity of the pre-trained student model.
Deep Mutual Learning (DML)
DML, also known as online or co-distillation, offers a perspective where two models, designated as student models $s_{1}$ and $s_{2}$, engage in a mutual learning process [14, 4, 15]. Unlike the traditional teacher-student dynamic in KD, DML involves a peer-to-peer collaborative learning approach:
$\mathcal{L}_{s_{1}}=\mathcal{L}_{CE}(p^{s_{1}},y)+\mathcal{L}_{KL}(p^{s_{1}},p^{s_{2}});$
$\mathcal{L}_{s_{2}}=\mathcal{L}_{CE}(p^{s_{2}},y)+\mathcal{L}_{KL}(p^{s_{2}},p^{s_{1}}).$ (3)
Here, losses $\mathcal{L}_{s_{1}}$ and $\mathcal{L}_{s_{2}}$ for student models $s_{1}$ and $s_{2}$ combine cross-entropy with true labels $y$ and KL divergence between each other’s probabilities.
Data Partitioning KD (DP-KD)
DP-KD [12] partitions the transfer data into two categories: one where the student model $s$ benefits from the teacher model $t$’s feedback, and another where the student’s pre-existing knowledge, represented by a frozen version of the initial student model (denoted as $st$), is preserved. Each data sample $l$ is assigned to one of these categories based on the highest prediction probability for the corresponding ground-truth class:
$m^{t}_{l}$ $=\textstyle\prod_{i}[p_{i}^{t}(z_{t,l})>p_{i}^{st}(z_{st,l})];$ (4)
$m^{st}_{l}$ $=\textstyle\prod_{i}[p_{i}^{t}(z_{t,l})\leq p_{i}^{st}(z_{st,l})].$
In this context, $m^{t}_{l}$ and $m^{st}_{l}$ are binary masks for each sample $l$, determining whether knowledge transfer from the teacher model $t$ is appropriate, or if the student model $s$ should retain its own knowledge. The distillation loss function is then defined as:
$\textstyle\mathcal{L}_{dist}=$ $\frac{T^{2}}{n}\sum_{l=0}^{n}\Big{[}m^{t}_{l}\mathcal{L}_{KL}\left(p^{s}(z_{s,l},T),p^{t}(z_{t,l},T)\right)$ (5)
$+m^{st}_{l}\mathcal{L}_{KL}\left(p^{s}(z_{s,l},T),p^{st}(z_{st,l},T)\right)\Big{]}.$
In the unsupervised variant of DP-KD, samples are assigned based on the maximum prediction probability for a sample, selecting the model with the highest confidence in its prediction. This model selection by confidence circumvents the need for true labels, enabling knowledge transfer to occur even without labeled data. We explore the efficacy of both the supervised and unsupervised paradigms of DP-KD to assess their respective impacts on the performance of pre-trained models.
II-C Related work
KD has seen notable expansion into distributed training, particularly in the transfer of knowledge among pre-trained models. This expansion has been increasingly prominent in fields such as distributed training and FL, as evidenced by a range of studies [4, 5, 9, 6, 11, 10, 7]. Techniques including FedMD [5], Cronus [11], and DS-FL [6] leverage KD for enhanced communication efficiency and to address model heterogeneity challenges. Decentralized frameworks like Def-KT[4] and ProxyFL [9] employ online distillation, facilitating collaborative learning in the absence of a central authority.
Despite these advancements, much of the research remains focused on singular KD strategies and a narrow range of data partitioning methods. Our study seeks to fill this gap by examining a variety of KD methodologies across different data distribution and transfer set scenarios, aiming to detail the performance nuances of diverse KD techniques in distributed training. This includes exploring the impact of hyperparameter tuning in KD, as highlighted by recent findings [13]. Furthermore, while existing FL studies have successfully integrated KD to reduce communication loads [6, 7, 8, 4], and support flexible model architectures across participants [5, 9, 10, 11], our study proposes a new approach to further cut communication overhead, particularly in cross-silo FL settings, enhancing FL system efficiency.
III Methodology
Based on the assumption of a set of participants (so-called clients), each with a private dataset, this study focuses on the application of KD techniques between pre-trained models, performed as follows:
  1. [itemsep=0pt, parsep=0pt]
  2. 1.
    Each model undergoes pre-training on its respective participant’s training dataset.
  3. 2.
    Knowledge is then transferred from the teacher to the student model utilizing one of the aforementioned KD techniques.
  4. 3.
    By default, the student model’s training dataset is the transfer set. We also examine the use of public datasets, both labeled and unlabeled, to evaluate their impact on the distillation process.
We assume a distributed training scenario with 10 participants for each data partitioning strategy. This setup results in 90 possible teacher-student pairings for each scenario, excluding self-transfers. We now describe the setup for this experimental methodology in detail.
Task, models, and datasets.  For this study, we focus on a standard computer vision task: image classification. As a dataset, we use CIFAR10 [16], which is commonly used in the FL literature (e.g., [17, 18, 19, 20, 21, 22]). The CIFAR-10 dataset consists of 32x32 color images of 10 classes. There are 50K training images and 10K test images. We use 10% of the training set as a validation set. CIFAR-10’s testing set is used as a shared dataset to measure the accuracy of the participants’ models before and after distillation.
To exclude the effects of different model architectures, we use ResNet-18 for each participant. We plan to revisit this choice in follow-up work.
For each participant, we pre-train the model over its training dataset for at most 100 epochs. We employ early stopping after the validation performance plateaus for 10 epochs. We use the Adam optimizer with a learning rate of $1\times 10^{-3}$, weight decay of $4\times 10^{-4}$, and a batch size of 32. These hyperparameters remain constant across all experiments.
Data partitioning strategies.  We experiment with the following five data partitioning strategies:
  1. 1.
    Uniform Data Distribution: Each participant $P_{i}$ possesses an equal number of data samples from all classes, resulting in balanced local datasets $D_{i}$, which promote uniform model training across the network.
  2. 2.
    Quantity Skew Distribution [23]: The size of each participant’s local dataset $|D_{i}|$ varies. A Dirichlet distribution, denoted as $q_{i}\sim\operatorname{Dir}(\beta)$, is used to model the size variability, where $q_{i}$ represents the proportion of data for participant $P_{i}$. The concentration parameter $\beta$, set to 0.5, controls the degree of data quantity skew among participants.
  3. 3.
    Specialized Non-IID Distribution: In this strategy, each participant $P_{i}$ primarily holds data from one class, making up 91% of their dataset. The remaining 9% is distributed evenly among samples from other classes, ensuring exposure to the entire class spectrum.
  4. 4.
    Label Skew Random Chunks Distribution: This approach involves assigning a random number of classes to each participant, followed by allocating random chunks, each containing a minimum of 100 samples, from these classes. This leads to a controlled variation in class distribution and dataset volume.
  5. 5.
    Label Skew Dirichlet Distribution: For each class $c$, we construct a distribution vector $p_{c}\sim\operatorname{Dir}_{K}(\bm{\beta})$, where $K$ is the number of participants and $\bm{\beta}$ is a vector of concentration parameters alternating between 0.1 and 0.5. The proportions $p_{c,k}$ derived from $p_{c}$ determine the allocation of class $c$’s samples to participant $P_{k}$, creating a skewed label distribution across participants, with lower $\beta$ values leading to greater skew.
KD setup.  We assess the effectiveness of three distinct KD methodologies: Vanilla KD, DML, and DP-KD. Each of these methodologies is explored with a trio of transfer set options:
  1. [itemsep=-0pt, parsep=0pt]
  2. 1.
    A public labeled dataset [12, 24, 25, 1].
  3. 2.
    A public unlabeled dataset [6, 11, 7, 20, 8, 26, 22, 19].
  4. 3.
    The student’s own training dataset [13, 4].
For Vanilla KD, the parameters used are $T=1$ and $\alpha=0.5$, reflecting common practices in existing literature [11, 7, 20, 8, 26, 4, 9]. When utilizing a public dataset as a transfer set, we incorporate a copy of the student’s pre-trained model as an additional teacher, based on recommendations from [13].
In the Tuned KD approach, we primarily focus on the optimization of the $\alpha$ parameter, followed by the tuning of the temperature parameter $T$. This optimization is conducted via a systematic grid search across the parameter space, as delineated in Table I.
For DML, we follow the methodology described in [4]. Specifically, from the two post-distillation models, we select the model whose dataset has been utilized as the default transfer set.
For DP-KD, we employ the unsupervised version when using a public unlabeled transfer set or the student’s dataset. Conversely, when using a public labeled transfer set, we opt for the supervised version of DP-KD, as outlined in Eq. 5.
TABLE I: Space of hyper-parameters.
Name Symbol Values
Temperature $T$ [0.1, 0.5, 1, 1.5, 2, 2.5, 3, 4, 5, 6, 7]
Weight $\alpha$ [0.1, 0.25, 0.5, 0.75, 0.9]
For simplicity and to align with our peer-to-peer collaborative system assumption, we predominantly focus on a single student and teacher setting. While we briefly delve into configurations with multiple teachers and a single student in § V, we acknowledge the complexity of these setups and reserve their extensive exploration for future research endeavors.
Metrics.  The primary metric of evaluation is the accuracy gain, defined as the change in percentage points of a participant’s model accuracy after KD compared to its pre-distillation accuracy. This gain may be negative, indicating a potential decline in performance due to KD.
Also, to obtain a balanced perspective on both the gains and potential drawbacks of KD, we assess forgetting, quantified as the decrease in per-class accuracy of the student model post-distillation relative to its pre-distillation performance, and learning, which is the increase in per-class accuracy of the student model post-distillation.
IV Multidimensional Evaluation of Knowledge Distillation Techniques
(a) Uniform Data Distribution
(b) Quantity Skew Distribution
(c) Specialized Non-IID Distribution
(d) Label Skew Random Chunks Distribution
(e) Label Skew Dirichlet Distribution
Figure 2: Evaluation of KD techniques across various data partitioning strategies and transfer set options.
We start with a comparison of various KD approaches, demonstrating their performance across different data distribution scenarios and transfer set configurations. Fig. 2 illustrates the results. Each subplot of the figure is a scatter plot showing knowledge transfer across all 90 participant pairs. That is, each data point in these subplots corresponds to a distinct knowledge transfer instance between two models. The x-axis denotes the performance differential between the teacher and the student models (quantified as teacher accuracy minus student accuracy), while the y-axis value reflects the accuracy gain achieved through KD.
Teachers that exhibit notably higher accuracy compared to their corresponding students are categorized as strong, while those with lower accuracy relative to the students are designated as weak. Instances of KD that result in improved accuracy are designated as positive knowledge transfers.
This comparative analysis includes transfer sets of varying sizes to determine their impact on the efficacy of KD. Specifically, the “Public unlabeled (large)” transfer set comprises 5,000 samples, the “Public unlabeled (small)” and the “Public labeled transfer set” each contains 500 samples, randomly sampled from the original dataset before partitioning. The “Student data as a transfer set” varies in size corresponding to the student dataset.
We now analyze our main findings from this comparison. We first contrast the performance of DML and DP-KD with Vanilla KD, which represents the baseline approach. Then, we illustrate the impact of hyperparameter tuning by comparing Tuned KD against other approaches.
IV-A Different transfer set options
Finding (1): Increasing the number of samples in the transfer set enhances the effectiveness of KD, leading to more pronounced positive knowledge transfers.
Our analysis reveals a proportional relationship between the size of the transfer set and KD effectiveness. An increase in the transfer set’s sample size consistently leads to enhanced outcomes of positive knowledge transfers, as illustrated in Fig. 2 (particularly evident when comparing positive knowledge transfers in columns 2 and 3). These findings underscore the importance of transfer set size in optimizing KD efficiency and offer practical insights for refining KD methodologies. Prior work [27] supports this observation, suggesting that an expanded transfer set can significantly reduce the student classifier’s loss, aligning the student’s parameters more closely with those of the teacher model.
It is worth noting that a larger sample size in the transfer set may also lead to more pronounced negative knowledge transfers under certain conditions, as observed with the DML approach in Fig. 2 (compare columns 2 and 3). As will be discussed next, the DML approach is not recommended in the absence of a student dataset or another public labeled transfer set.
Finding (2): The DML approach exhibits poor performance in unsupervised transfer set settings.
Our findings indicate that while DML generally outperforms Vanilla KD and DP-KD in scenarios involving weaker teachers, particularly with ‘Student data as a transfer set’ and ‘Public labeled transfer set’, its performance significantly diminishes in the absence of labeled data. In such cases, DML often results in negative knowledge transfers, as observed in columns 2 and 3 of Fig. 2. This pattern cautions against employing DML when neither student data nor a labeled transfer set is accessible.
IV-B DML and DP-KD vs. Vanilla KD
Finding (3): Vanilla KD demonstrates similar or superior performance compared to DP-KD across various data distribution scenarios.
Fig. 2 provides a detailed comparative analysis underscoring this finding. In scenarios such as Uniform Data Distribution, Quantity Skew Distribution, and Label Skew Random Chunks Distribution, Vanilla KD’s performance is comparable to that of DP-KD. However, it notably excels in the Label Skew Dirichlet Distribution scenario and demonstrates significant superiority in the Specialized Non-IID Distribution context, across all tested transfer set options. This pattern of performance highlights the robustness and adaptability of Vanilla KD, particularly in complex data distribution scenarios and when paired with teachers of comparable or superior strength.
Finding (4): DP-KD performs poorly in the ‘Specialized Non-IID Distribution’ scenario.
This outcome suggests that the model’s guidance may be less effective in scenarios characterized by significant class imbalances—even though the teacher model has been exposed to samples from all classes. Such a scenario can lead to misleading confidence from the teacher model, ultimately resulting in suboptimal knowledge transfer and poor performance outcomes. This underscores the importance of considering class balance in the teacher’s training dataset for effective DP-KD application.
Finding (5): There is a notable susceptibility in employing Vanilla KD and DP-KD methods with weak teachers, especially when using student data as the transfer set.
In contrast to Vanilla KD and DP-KD, DML exhibits a distinctive ability to facilitate positive knowledge transfers with weaker teachers when utilizing student data as the transfer set. This is clearly demonstrated in Fig. 2, particularly in column 1. Such findings highlight DML’s potential in scenarios where the teacher model is less robust, emphasizing its effectiveness in optimally leveraging student data. However, in scenarios involving stronger teachers, Vanilla KD tends to outperform DML.
(a) Uniform Data Distribution
(b) Quantity Skew Distribution
(c) Specialized Non-IID Distribution
(d) Label Skew Random Chunks Distribution
(e) Label Skew Dirichlet Distribution
Figure 3: Optimization of hyperparameters $T$ and $\alpha$ in KD across various data partitioning strategies. The first column illustrates optimal $T$ values from grid search, related to performance differences between student and teacher models. The second column shows the best $\alpha$ values. The third column presents cumulative accuracy gains in percentage points from different KD methods. The final column details the accuracy gains for each participant pair using various KD techniques.
(a) Label Skew Random Chunks Distribution, CINIC10
Figure 4: Optimization of hyperparameters $T$ and $\alpha$ in KD for CINIC-10 dataset.
IV-C Optimizing hyperparameters: Tuned KD effectiveness
The optimization of hyperparameters is crucial in enhancing the efficiency of KD, a notion supported by recent research [13]. This subsection focuses on the efficacy of the Tuned KD approach, specifically targeting the optimization of the weight $\alpha$ and then the temperature $T$. This optimization process is conducted via a grid search over the parameter space outlined in Table I. For this analysis, and in the subsequent subsections unless otherwise noted, we use student data as the transfer set.
Fig. 3 illustrates the main outcomes of our hyperparameter optimization endeavors. It highlights the impact of varying $T$ and $\alpha$ values on KD performance across different data partitioning strategies and identifies the optimal settings for these parameters.
Figure 5: Comparative analysis of the best teachers across diverse KD scenarios. The bar charts in the first row depict the frequency of each participant being selected as the best teacher within four different KD approaches, across five data distribution strategies. The heatmaps in the second row illustrate the corresponding participant accuracies, providing insight into the relationship between a participant’s accuracy and their likelihood of being an optimal teacher. The x-axis of the first row represents the participant’ IDs, which align directly above their respective accuracy heatmap.
Finding (6): Tuned KD outperforms all the other KD approaches across all data partitioning strategies.
As demonstrated in the last column of Fig. 3, Tuned KD consistently outperforms Vanilla KD, DML, and DP-KD in all data distribution scenarios. It is particularly noteworthy that Tuned KD achieves positive knowledge transfers even with weak teachers, markedly differing from the outcomes seen with Vanilla KD and DP-KD. Additionally, in situations involving strong teachers, Tuned KD enhances accuracy gains, surpassing all other KD approaches. Furthermore, the performance analysis summarized in the third column of Fig. 3 depicts the cumulative accuracy gain achieved by each KD method across different data distributions, reinforcing the standout performance of Tuned KD. This finding underscores the importance of optimally tuning the $T$ and $\alpha$ parameters. It also opens up new avenues for future research, particularly in developing methods for efficient and automated tuning of these parameters tailored to the specific setting and learning objectives.
Finding (7): Tuning $T$ and $\alpha$ parameters becomes particularly important when the student and the teacher have diverse performances, while it becomes less significant when they have comparable performances.
The last column of Fig. 3 illustrates this phenomenon clearly. In scenarios where there is a significant performance disparity between the teacher and the student, Tuned KD demonstrates a marked improvement over Vanilla KD. Conversely, in cases where the teacher and student exhibit similar performance levels, the benefits of tuning $T$ and $\alpha$ diminish, resulting in Tuned KD achieving performance levels comparable to those of Vanilla KD.
Finding (8): Insights about the best $T$ and $\alpha$ values.
Fig. 3, particularly columns 1 and 2, illustrates the optimal $T$ and $\alpha$ values determined through our grid search across diverse data distributions. While identifying a consistent pattern for the best $T$ and $\alpha$ values in the Uniform Data Distribution is challenging, clear trends emerge in the Quantity Skew Distribution, the Label Skew Random Chunks Distribution, and the Label Skew Dirichlet Distribution. In these scenarios, a smaller $T$ value tends to be more effective with a weaker teacher, while a larger $T$ value is advantageous for a stronger teacher. Correspondingly, the ideal $\alpha$ value, indicating the weight assigned to the teacher’s output, should align with the teacher’s relative strength compared to the student. Thus, a significantly stronger teacher warrants a higher $\alpha$ value (above 0.5), while a weaker teacher should receive a lower weight. Moreover, the Specialized Non-IID Distribution generally favors lower values of $T$ and $\alpha$, leading to more favorable outcomes across all teacher-student pairings.
Fig. 4 includes experiments using the CINIC10 dataset [28], albeit with a singular data distribution approach, serves to validate and reinforce our findings from the CIFAR10 dataset. This consistency indicates result robustness and suggests the potential generalizability of our conclusions across different datasets, opening avenues for further exploration in diverse data environments.
IV-D Best teachers across KD methods and data distributions
Fig. 5 juxtaposes the frequency of participants emerging as the best teachers against their respective performance within varied data distribution strategies, utilizing the four KD approaches. The best teacher is defined as one that provides the greatest accuracy gain for a given student participant. This analytical approach aims to discern the potential correlation between a participant’s accuracy and its propensity to be an effective teacher.
Figure 6: Knowledge distillation techniques against varied data distribution characteristics.
Finding (9): Participant’s accuracy is not the sole determinant of its efficacy as a teacher in KD.
The findings indicate that participant accuracy, though pivotal, is not the only factor influencing the selection of the best teacher across diverse KD techniques and data distributions. This highlights the complexity of the KD teacher selection process, which appears to consider a broader spectrum of factors.
Under the uniform data distribution, the emergence of the best teachers appears to be evenly spread across participants for most KD approaches. In contrast, in scenarios with significant data skewness, particularly in quantity and label skew distributions, KD methods such as Tuned KD, Vanilla KD, and DP-KD tend to prefer teachers with high accuracy, but also possibly incorporate other factors in their selection.
Interestingly, the DML approach’s teacher selection exhibits a weaker correlation with accuracy, implying that it may value additional attributes such as the diversity or specificity of participant knowledge.
This intricate selection dynamic indicates an opportunity to enhance KD algorithms by integrating a more complex array of criteria beyond accuracy. Subsequent studies should aim to decode these implicit factors, potentially leading to more refined and efficacious KD paradigms.
(a) Uniform Data Distribution
(b) Quantity Skew Distribution
(c) Specialized Non-IID Distribution
(d) Label Skew Random Chunks Distribution
(e) Label Skew Dirichlet Distribution
Figure 7: Comparative analysis of learning vs. forgetting in KD across diverse data distributions.
Figure 8: Decision tree to determine the suitable KD approach.
IV-E Revisiting KD methods: A comparative perspective through data distributions and learning dynamics
In this subsection, we re-examine the performance of various KD methods, including Tuned KD, under a different light. This revisit emphasizes a comparative analysis grounded in different data distribution strategies, consolidating our earlier observations and unveiling new insights into the adaptability and robustness of these methods.
Fig. 6 juxtaposes the efficacy of KD methods against varied data distributions. This comparative visualization allows to discern patterns and outliers in performance, highlighting the consistent strength of Tuned KD across multiple scenarios. By contrasting these methods side-by-side, we gain a clearer understanding of how each method responds to diverse data conditions.
Fig. 7 investigates the learning-forgetting (cf. § III) trade-off in KD methods. Tuned KD and DML show less forgetting with weaker teachers, outperforming Vanilla KD and DP-KD. Uniform data distributions reveal similar performances across methods, with a minor uptick in forgetting for DP-KD. Tuned KD emerges as the most effective, maximizing learning in all scenarios and reducing forgetting in Quantity Skew and Specialized Non-IID distributions. In contrast, it shows increased forgetting within Label Skew distributions. DML shows minimal forgetting but is constrained in learning gains, suggesting a conservative knowledge consolidation preference. This analysis underscores the need to tailor KD approach selection to the specific strengths and weaknesses of each method, aligned with data characteristics and learning goals.
Finally, the decision tree in Fig. 8 serves as a strategic guide for selecting suitable KD methods based on specific use cases and data environments, offering a practical approach for applying KD methods in various contexts.
V KD from Multiple Pre-trained Models
This section explores the application of KD in the context of multiple pre-trained models. Our investigation primarily revolves around Vanilla KD, examining its effectiveness in consolidating knowledge from various models, each with distinct training experiences and data exposures. This study aligns with the broader goal of harnessing the collective intelligence of diverse models in decentralized and federated learning systems. Additionally, the section briefly touches upon the application of KD in enhancing FL, assessing its role in pre-consolidating models before initiating the FL process. This aspect of the study emphasizes the potential of KD in streamlining and advancing the FL workflow.
For FL experiments, we follow the optimized Federated Averaging (FedAvg) parameters from [29], chosen for their relevance to our CIFAR-10 dataset and ResNet-18 model setup. We use an SGD solver with a learning rate of 0.01, weight decay of $4\times 10^{-4}$, no momentum, and local epochs set to 2. We implement a standard cross-silo FL setting with full participation (participation rate of 1) across 100 communication rounds.
TABLE II: Combining multiple models across various data distributions and transfer set options
Data Distribution
Public Unlabeled Transfer Set Student Data as Transfer Set
Worst-Student Best-Student Untrained-Model Worst-Student Best-Student
Weighted Unweighted Weighted Unweighted Weighted Unweighted Weighted Unweighted Weighted Unweighted
Uniform
71.64 72.36 72.35 73.43 69.55 70.59 71.34 72.24 72.68 73.97
Quantity Skew
68.18 64.89 77.90 73.86 67.31 64.66 29.77 30.01 78.34 76.33
Specialized Non-IID
51.57 54.72 53.07 55.11 52.55 52.98 43.70 50.72 52.39 53.60
Label Skew Random Chunks
65.44 66.73 73.33 71.68 66.35 66.22 19.22 20.67 75.71 75.41
Label Skew Dirichlet
61.67 54.03 67.63 57.82 63.77 52.41 28.61 27.35 66.92 63.00
V-A Combining multiple models
We delve into methodologies for the effective amalgamation of knowledge across multiple pre-trained models via KD. Our focus includes critical aspects such as the selection of an appropriate starting student model and the assessment of various teacher weighting strategies across different transfer set scenarios.
Where to start?  Selecting the optimal starting student model is key in KD. We explore three scenarios: using the lowest-performing, highest-performing, or an untrained student model.
Transfer set.  Our approach incorporates two transfer set options: a public unlabeled dataset with 5,000 samples, and the student’s dataset. The models undergo 30 epochs of training over these transfer sets.
Teachers’ weights.  We investigate equal weighting for all teachers and adaptive weighting, which is based on each teacher’s class-wise contribution.
In adaptive weighting, the weight vector $\alpha_{j}$ for each teacher $j$ considers their accuracy $Acc_{j}^{c}$ for each class $c$, relative to the student’s accuracy $Acc_{s}^{c}$ and the total accuracy of all teachers for that class:
$\alpha_{j}^{c}=\begin{cases}\frac{Acc_{j}^{c}}{Acc_{s}^{c}+\sum_{i=1}^{J}Acc_{i}^{c}},&\text{if }Acc_{s}^{c}+\sum_{i=1}^{J}Acc_{i}^{c}>0\\ 0,&\text{otherwise}\end{cases}$ (6)
Here, $J$ is the total number of teachers, and the dimension of each $\alpha_{j}$ vector matches the number of classes.
The $\mathcal{L}_{KL}$ in Eq. 2 can be updated considering the combined influence of multiple teachers, as:
$\mathcal{L}_{KL_{multi}}=\sum_{j=1}^{J}\alpha_{j}\>\mathcal{L}_{KL}(p^{s},p^{j}).$ (7)
This equation integrates the KL divergence losses between the student’s predictions and each teacher’s predictions, each scaled by its respective $\alpha_{j}$.
Table II presents a summary of our experimental outcomes, offering insights into the efficacy of different KD strategies under varying conditions. The ‘Weighted’ and the ‘Unweighted’ columns in the table represent the adaptive weighting and the equal weighting strategies, respectively.
Finding (10): Selecting the highest-performing model as the student in the consolidation process consistently yields the most favorable outcomes.
The data suggest a consistent trend: initiating the distillation process with the best-performing student model generally leads to better results. This could be attributed to the more advanced starting point it provides, requiring less knowledge to assimilate.
Finding (11): The adaptive weighting approach shows varied effectiveness based on data distribution, with marked gains in Label Skew Dirichlet and Quantity Skew scenarios.
The adaptive weighting approach delineated in Eq. 6 and Eq. 7 manifests considerable enhancements in specific data distribution contexts, notably the Label Skew Dirichlet and Quantity Skew scenarios. Conversely, in the Specialized Non-IID and the Uniform data distributions, the equal weighting strategy across all teachers appears to be more beneficial. This variability underscores the need to tailor weighting strategies based on specific data contexts in KD applications.
Moreover, the choice between utilizing the best student data or a public unlabeled dataset as the transfer set did not show significant differences in performance, suggesting flexibility in transfer set selection based on availability.
Figure 9: Comparative performance of traditional FedAvg vs. KD Pre-consolidated FedAvg across various data partitioning strategies.
However, the comparison between using the best student’s data and the worst student’s data as the transfer set shows a marked difference. Opting for the best student’s data typically leads to significantly better results. This can be attributed to the broader and possibly richer data representation in the best student’s dataset, in contrast to the more limited scope of the worst student’s data. This finding highlights the importance of transfer set size and quality, corroborating the initial observations we made in § IV regarding the impact of transfer set on KD performance.
In light of these observations, we will proceed with further experiments using the student data as the transfer set and employing the adaptive weighting mechanism for knowledge distillation.
V-B Accelerated FL with KD pre-consolidation
This subsection explores the application of KD to enhance FL. We assess the effectiveness of using KD for model pre-consolidation before initiating FL, particularly in cross-silo settings, a domain directly relevant to our research.
We compare traditional FL using FedAvg with the proposed approach where models first undergo KD pre-consolidation. This method utilizes the pre-consolidated model as the initial model for FL, diverging from the standard practice of starting with randomly initialized models. Our analysis spans various data distribution scenarios to gauge the impact of KD pre-consolidation on reducing the number of communication rounds necessary to achieve target accuracy in FL.
Fig. 9 presents the results of this comparative analysis.
Finding (12): Employing KD for model pre-consolidation as an initial step in FL can significantly accelerate the learning process.
KD consolidation, when used alone, might not always produce satisfactory results in complex environments like Specialized Non-IID and Label Skew Dirichlet distributions. However, incorporating it as a preliminary step in FL significantly speeds up the learning process. This demonstrates the potential of KD pre-consolidation as a powerful tool for improving the efficiency of FL systems, particularly in accelerating the convergence towards target accuracy levels.
VI Conclusion
This paper presents an in-depth exploration of KD in the context of pre-trained models, offering vital insights and empirical evidence into its effectiveness. We have systematically compared various KD methodologies, such as Vanilla KD, Tuned KD, DML, and DP-KD, across multiple data distribution strategies, highlighting their unique strengths and applications. Our findings demonstrate that Tuned KD, particularly with optimized parameters, consistently outperforms other methods in various scenarios.
Furthermore, we have unveiled the potential of KD in accelerating FL. By pre-consolidating models using KD, our experiments indicate a significant reduction in the number of communication rounds required for FL, thus enhancing its efficiency.
Overall, our research enriches the understanding of KD within distributed learning systems. It underscores KD’s strategic value in enhancing both model performance and the efficiency of training processes, setting the stage for further advancements in both centralized and decentralized distributed training environments.